2 0 If A is a square matrix then A2 = AA. Let A = Find A2 3 1 A2 = (Simplify your answer.)
Answer #1

\small \text{Given, } A = \begin{bmatrix} 2 & 0\\ 3 & 1 \end{bmatrix} \\ \begin{align*} A^2 = AA &= \begin{bmatrix} 2 & 0\\ 3 & 1 \end{bmatrix} \begin{bmatrix} 2 & 0\\ 3 & 1 \end{bmatrix} \\ &= \begin{bmatrix} (2*2 + 0*3) & (2*0 + 0*1)\\ (3*2 + 1*3) & (3*0 + 1*1) \end{bmatrix} \\ &= \begin{bmatrix} 4 & 0\\ 9 & 1 \end{bmatrix} \end{align*}